OpenStudy (anonymous):

Equation 1: x=(*inital* cos *theta*)z Equation 2: y=(*inital* sin *theta*) z - 16(z^2) Solve for z in the first equation and then subsitute the expression for z into the second equation. Simplify so there are no trig functions in any denominator in the equation for y. You now should have an equation which shows y as a function of x. Okay~ I got most of the answer, but I'm confused as to how MUCH I should simplify. And later i have to graph this with theta being 45 degrees and inital being 100. Oh, just for the sake of ranting. I am on break with a ton of hw. happy holidays!

OpenStudy (underhill):

Rainbow, There are two conditions for the simplified problem here: (1) y alone on one side of the equation and (2) "no trig functions in any denominator." The following then, would be my simplification: \[y = x \tan \theta - 16x^2\sec^2\theta\]

OpenStudy (underhill):

If theta is constant, this will graph to be an inverted parabola.

OpenStudy (underhill):

What does Rainbow mean by "initial?"

OpenStudy (anonymous):

x=z cos t, y=z sin t - 16z^2, z=x/cos t, sub into y gives y=x tan t - 16 x^2 sec^2 t you can plot or graph these